OpenStudy (anonymous):

Help.. Ok 4 out of 22 students walk to school.if two of the 22 students are selected at random,what is the probability that both walk to school

OpenStudy (anonymous):

\[4/22 * 3/21\]

OpenStudy (anonymous):

you take the probability that you get a student that walks to school on the first selection\[\frac{ 4 }{ 22 }\]and multiply that by the probability that you get a student that walks on the second, remembering that you already picked one of those students out, so you have that the probability of the second student walking is\[\frac{ 3 }{ 21 }\]by multiplying these together you will get the probability that both of them happen, so\[\frac{ 4 }{ 22 }x \frac{ 3 }{ 21 }=\frac{ 12 }{ 462 }\] and you can go ahead and simplify that further if you need

OpenStudy (anonymous):

Thanks for your help guys now I finished my worksheet